### v1

@@ 4.1 Common Classifier Head
Typically the classifier head of a transformer-based model used for the AES consists of a simple dense layer or conjunction of dense layers and LSTMs (or Bi-LSTMs) [44] [46] [21], so we decided that this design would represent performance as a standard design choice and the graded decrease in size in the FFNs would provide the model with a bit of stability.
To counteract the high correlation as found out by Jeon and Strube [15], we use prompt-specific normalisation discussed in detail in Section 5.2.2.
@@ 4.2 Baseline
The next question we must ask is —  how would the baseline perform without any added context? To answer this, we look into establishing a baseline with a transformer model. The architecture used is RoBERTa with the common classifier head. The results of this experiment can be seen in Table 6.1. In contrast to our design, most of the literature that uses transformer-based models have used BERT [44] [46] [21], and some used Longformer[3] as well [44]. The classifier head for most of these implementations consisted of a simple, densely connected linear layer or a combination of LSTMs and dense layers. Although we use a standard architecture, our baseline’s primary purpose is not to compare model implementations from literature but to provide a baseline score for later model performances. A point to note here is that since we use RoBERTa instead of BERT, we expect our baseline to outperform those in the preexisting literature.
@@ 4.3 Input Length
This made us reconsider our approach and choose Longformer [3] for our transformer architecture. Longformer is based on RoBERTa, and we must note that RoBERTa does outperform DeBERTaV3. The modifications made to the Longformer architecture allow it to handle longer documents much better than RoBERTa. Thus, another baseline was established for Longformer using only raw data input. Table 6.1 shows that the baseline Longformer model outperformed all others.
@@ 4.4 Relative context: Margin Ranking
Now that we have established that untruncated input would be ideal for transformer-based models, it is time for us to start looking at context augmentation, starting with relative context. To allow the model to understand the relative ordering of essays, we introduce Margin Ranking (MR) loss. MR loss was proposed by Wang et al. [44] to help the model understand the correct order of essays in a batch. The order of essays is crucial for scoring, as allowing the model to capture this aspect encourages the model to penalise incorrect ordering. For each batch of essays, we look at every possible pair of essays and calculate it as follows:
@@ 4.6.1 EDU Span Prediction
Before using EDU-labelled data, we develop an EDU span classification model based on Wang et al.[43]. The reason we chose their work to base our architecture on is because of its high adaptation rate as a standard model to do the upstream task of EDU labelling [28] [45] [7] [6][37]. It is also open-source, allowing us to change the architecture as our needs are. The standard model also has quite an impressive F1-Score of 0.940. Figure 4.3 describes the model architecture graphically. Our variation implements RoBERTa embeddings with the original BiLSTM-CRF network architecture. RoBERTa was our choice of text embedding generator as it has a much better understanding of language than the original 300-D Glove embeddings. Also, there were no truncation concerns since it was used to get sentence-level embeddings and not document or essay-level. BiLSTMs address context loss in longer sequences. Long-range dependencies, crucial for EDU boundaries, are managed using Restricted Self Attention (RSA), discussed in Section 4.6.1. This was arguably the most essential contribution of Wang et al. in the original paper. The Conditional Random Field used with the Bi-LSTM also significantly contributes to the performance superiority and is discussed in detail below.
@@ 4.7.1 AC Span Prediction
The EDU span boundaries were used under advisement from the works of Peldszus and Stede [33] [34], as also Musi et al. [27] who demonstrated a correlation between discourse relations from Rhetorical Structure Theory (RST) and argumentative relations by translating the discourse structure mapping into argumentation structures.
@@ 4.8.2 Argument Component with Prompts and Feature Context
In this section, we discuss exploiting other feature-based contexts from the dataset. For instance, we would like to refer to the discussion of the work of Jeon and Strube [15] from Section 2.2. They pointed out that the resolved score in the dataset is heavily correlated to the lengths of the input essays.
@@ 5.1 Evaluation Metrics
As for the ASAP-AES dataset, we used the Quadratic Weighted Kappa (QWK) score to benchmark our model’s performance against others in the available literature. This is partially because this was the benchmarking tool advised by the ASAP contest on Kaggle and has been used as a benchmarking method for most if not all, papers on this dataset. However, elaborating further, the AES model aims to correlate scores predicted by the model with the human scores from the dataset. QWK scores consider the model’s ability to predict scores and reflect its consistency with human judgments. The essays in the dataset are scored on an ordinal scale, where the scores have a natural order but may not have a fixed interval between them. QWK handles ordinal data very well [9] and assesses the agreement between different raters or systems based on these ordered categories. Another thing to note is that the QWK is an extension of Cohen’s Kappa metric, commonly used to enumerate inter-rater agreements. But as opposed to Cohen’s Kappa metric, QWK assigns different weights to different levels of agreement/disagreement based on how far apart the scores are. This is particularly useful for our use case, where the distance between scores matters more than agreement or disagreement. It is also robust to chance agreements between raters. This is particularly useful as multiple human raters might score the same essay differently due to various factors. QWK adjusts for the possibility of chance agreement and provides a more accurate assessment of model performance. Like in many real-world scenarios, the ASAP dataset has non-linear scoring scales. The difference between a score of 3 and 4 might have a different impact than the difference between 7 and 8. QWK is better suited to capture these nuances.
@@ 5.2.1 EDU and AC Span Labelling
The same parameters were used to train the model architecture used in EDU and AC labelling. The RST Discourse Treebank provides segmented EDU spans separately, whereas the AAE V2 dataset directly gives us the character-level spans of the ACs. These generated IO (Inside Outside) labels for the two tasks from the data. Negative Log Likelihood loss was used to train the models (for EDU and AC separately) along with L2 regularisation, and Adam [17] was used for the optimiser. F1-score was used as the final metric to judge the efficacy of these models. The model architecture used is as described in Figure 4.3.
@@
Adam [17] (Adaptive Moment Estimation) is an optimisation algorithm. It combines the advantages of both Adagrad [10] (which adapts learning rates based on historical gradient information) and RMSProp [41] (which uses moving averages of squared gradients) to provide fast and effective optimisation.
@@ 5.2.2 Automated Essay Scoring
In the Automated Essay Scoring task, just Mean Squared Error loss was initially used, but later, the combined weighted loss of MSE and Margin Ranking (MR) was implemented. RMSProp[41] was used as the optimising algorithm, and scikitlearnscikit-learn’s StandardScalerStandardScaler was used to normalise the scores. The early stoppage technique was used to save computation time if the validation QWK score did not improve in 15 epochs, and validation QWK scores were used to keep the best-performing model.
@@
RMSProp [41] was primarily chosen for the AES model for its stable and reliable convergence behaviour. Its gradual adjustments to learning rates make it particularly suited for non-convex optimisation problems, where finding the best solution is challenging. Unlike algorithms like Adam, which can lead to rapid changes in learning rates, RMSProp’s approach avoids potential instability during training, ensuring a smoother learning process.
@@ 5.3 Competing Models
    DeLAES Model: DeLAES, developed by Tashu et al. [39], is a novel approach that employs an n-gram-inspired CNN-Bi-GRU architecture. This model has achieved remarkable success, exhibiting a state-of-the-art mean QWK score of 0.903 across various benchmark datasets. Combining convolutional neural networks (CNNs) and bidirectional gated recurrent units (Bi-GRUs), their innovative strategy to capture local and contextual information in essays for enhanced grading accuracy along the granularities of bi-grams, tri-grams and 4-grams.
@@
    Tran-BERT-MS-ML-R Model: Wang et al. [44] proposed the Tran-BERT-MS-ML-R model, which leverages the power of BERT-based transformer architectures. This model takes a multi-faceted approach by incorporating transfer learning techniques and specialised methodologies to improve the performance of BERT models. By addressing specific challenges associated with essay scoring, Wang et al. achieved notable advancements in scoring accuracy.
@@
    $R^{2}$ BERT Model: $R^{2}$ BERT, introduced by Yang et al. [46], has been the state-of-the-art for transformer-based models for AES for a long time. The name $R^{2}$ BERT stems from its utilisation of both Regression and Ranking techniques for score prediction. This unique combination enhances the model’s ability to capture intricate nuances in essays, improving grading performance. Yang et al. demonstrated the model’s efficacy through comprehensive experimentation and comparisons with existing approaches.
@@
We began by investigating the impact of truncation, a critical consideration for models utilising BERT-based architectures with token limitations. We enhanced the model’s understanding by preserving the direct context of essays through non-truncation. Following this, we further delved into providing a relativistic ordering context via margin ranking, as detailed in Section 4.4. This entailed furnishing the model with a sense of the batch’s essay score ordering to enhance its evaluation capabilities. Further, we examined the augmentation of essays with their respective prompts. This approach improved the model’s grasp of essay quality and its comprehension of nuanced grading patterns within each essay set. We further explored context granularity through the levels of Elementary Discourse Unit (EDU) and Argument Component (AC). Inspired by Wang et al.’s work [43], we devised models for the upstream tasks of EDU and AC span labelling. These labelled essays formed the input for our transformer-based architecture, where we observed performance improvements.
@@ 6.2 EDU Span Classification
As discussed before, we enhanced the transformer’s grasp of individual text segments and their relationships by leveraging the EDUs. To get EDU labels, we implemented an EDU span classification model, discussed here. Our implementations were based on Wang et al.’s work [43], and we adapted their architecture to use RoBERTa embeddings instead of Glove and ElMo embeddings.
@@ 6.5 Comparing Models
Tran-BERT-MS-ML-R [44] 0.834 0.716 0.714 0.812 0.813 0.836 0.839 0.766 0.791
$R^{2}$ BERT [46] 0.817 0.719 0.698 0.845 0.841 0.847 0.839 0.744 0.794
@@
  • Graph Convolutional Networks (GCNs): Inspired by the work of Ruiz et al. [36], we may even want to use GCNs to exploit the argument graph structure for further context.
  • Finetuning: Finetuning the transformer-based models is also challenging, as explained in the paper by Wang et al. [44]. We can explore back-translation and relativistic scoring to augment data for further training instances to stabilise fine-tuning as explored by Jong et al. [16].
An objective study of these would only positively benefit the literature available in AES. It is important to note that the end-to-end argument graph mining models need to achieve better SOTA results [37] to avoid injecting noise into the transformer models instead of enriching them.

### v2

@@ 4.1 Common Classifier Head
Typically the classifier head of a transformer-based model used for the AES consists of a simple dense layer or conjunction of dense layers and LSTMs (or Bi-LSTMs) [wang_2022_on] [yang-etal-2020-enhancing] [liang2021rdrop], so we decided that this design would represent performance as a standard design choice and the graded decrease in size in the FFNs would provide the model with a bit of stability.
To counteract the high correlation as found out by Jeon and Strube [jeon-strube-2021-countering], we use prompt-specific normalisation discussed in detail in Section 5.2.2.
@@ 4.2 Baseline
The next question we must ask is —  how would the baseline perform without any added context? To answer this, we look into establishing a baseline with a transformer model. The architecture used is RoBERTa with the common classifier head. The results of this experiment can be seen in Table 6.1. In contrast to our design, most of the literature that uses transformer-based models have used BERT [wang_2022_on] [yang-etal-2020-enhancing] [liang2021rdrop], and some used Longformer[beltagy2020longformer] as well [wang_2022_on]. The classifier head for most of these implementations consisted of a simple, densely connected linear layer or a combination of LSTMs and dense layers. Although we use a standard architecture, our baseline’s primary purpose is not to compare model implementations from literature but to provide a baseline score for later model performances. A point to note here is that since we use RoBERTa instead of BERT, we expect our baseline to outperform those in the preexisting literature.
@@ 4.3 Input Length
This made us reconsider our approach and choose Longformer [beltagy2020longformer] for our transformer architecture. Longformer is based on RoBERTa, and we must note that RoBERTa does outperform DeBERTaV3. The modifications made to the Longformer architecture allow it to handle longer documents much better than RoBERTa. Thus, another baseline was established for Longformer using only raw data input. Table 6.1 shows that the baseline Longformer model outperformed all others.
@@ 4.4 Relative context: Margin Ranking
Now that we have established that untruncated input would be ideal for transformer-based models, it is time for us to start looking at context augmentation, starting with relative context. To allow the model to understand the relative ordering of essays, we introduce Margin Ranking (MR) loss. MR loss was proposed by Wang et al. [wang_2022_on] to help the model understand the correct order of essays in a batch. The order of essays is crucial for scoring, as allowing the model to capture this aspect encourages the model to penalise incorrect ordering. For each batch of essays, we look at every possible pair of essays and calculate it as follows:
@@ 4.6.1 EDU Span Prediction
Before using EDU-labelled data, we develop an EDU span classification model based on Wang et al.[wang-etal-2018-toward]. The reason we chose their work to base our architecture on is because of its high adaptation rate as a standard model to do the upstream task of EDU labelling [ni-etal-2019-justifying] [xu2020discourseaware] [chen2020modeling] [chakrabarty2020ampersand][saha-etal-2022-edu]. It is also open-source, allowing us to change the architecture as our needs are. The standard model also has quite an impressive F1-Score of 0.940. Figure 4.3 describes the model architecture graphically. Our variation implements RoBERTa embeddings with the original BiLSTM-CRF network architecture. RoBERTa was our choice of text embedding generator as it has a much better understanding of language than the original 300-D Glove embeddings. Also, there were no truncation concerns since it was used to get sentence-level embeddings and not document or essay-level. BiLSTMs address context loss in longer sequences. Long-range dependencies, crucial for EDU boundaries, are managed using Restricted Self Attention (RSA), discussed in Section 4.6.1. This was arguably the most essential contribution of Wang et al. in the original paper. The Conditional Random Field used with the Bi-LSTM also significantly contributes to the performance superiority and is discussed in detail below.
@@ 4.7.1 AC Span Prediction
The EDU span boundaries were used under advisement from the works of Peldszus and Stede [Peldszus2015AnAC] [peldszus-stede-2016-rhetorical], as also Musi et al. [musi-etal-2018-multi] who demonstrated a correlation between discourse relations from Rhetorical Structure Theory (RST) and argumentative relations by translating the discourse structure mapping into argumentation structures.
@@ 4.8.2 Argument Component with Prompts and Feature Context
In this section, we discuss exploiting other feature-based contexts from the dataset. For instance, we would like to refer to the discussion of the work of Jeon and Strube [jeon-strube-2021-countering] from Section 2.2. They pointed out that the resolved score in the dataset is heavily correlated to the lengths of the input essays.
@@ 5.1 Evaluation Metrics
As for the ASAP-AES dataset, we used the Quadratic Weighted Kappa (QWK) score to benchmark our model’s performance against others in the available literature. This is partially because this was the benchmarking tool advised by the ASAP contest on Kaggle and has been used as a benchmarking method for most if not all, papers on this dataset. However, elaborating further, the AES model aims to correlate scores predicted by the model with the human scores from the dataset. QWK scores consider the model’s ability to predict scores and reflect its consistency with human judgments. The essays in the dataset are scored on an ordinal scale, where the scores have a natural order but may not have a fixed interval between them. QWK handles ordinal data very well [DELATORRE2018144] and assesses the agreement between different raters or systems based on these ordered categories. Another thing to note is that the QWK is an extension of Cohen’s Kappa metric, commonly used to enumerate inter-rater agreements. But as opposed to Cohen’s Kappa metric, QWK assigns different weights to different levels of agreement/disagreement based on how far apart the scores are. This is particularly useful for our use case, where the distance between scores matters more than agreement or disagreement. It is also robust to chance agreements between raters. This is particularly useful as multiple human raters might score the same essay differently due to various factors. QWK adjusts for the possibility of chance agreement and provides a more accurate assessment of model performance. Like in many real-world scenarios, the ASAP dataset has non-linear scoring scales. The difference between a score of 3 and 4 might have a different impact than the difference between 7 and 8. QWK is better suited to capture these nuances.
@@ 5.2.1 EDU and AC Span Labelling
The same parameters were used to train the model architecture used in EDU and AC labelling. The RST Discourse Treebank provides segmented EDU spans separately, whereas the AAE V2 dataset directly gives us the character-level spans of the ACs. These generated IO (Inside Outside) labels for the two tasks from the data. Negative Log Likelihood loss was used to train the models (for EDU and AC separately) along with L2 regularisation, and Adam [kingma2017adam] was used for the optimiser. F1-score was used as the final metric to judge the efficacy of these models. The model architecture used is as described in Figure 4.3.
@@
Adam [kingma2017adam] (Adaptive Moment Estimation) is an optimisation algorithm. It combines the advantages of both Adagrad [adagrad] (which adapts learning rates based on historical gradient information) and RMSProp [tieleman2012lecture] (which uses moving averages of squared gradients) to provide fast and effective optimisation.
@@ 5.2.2 Automated Essay Scoring
In the Automated Essay Scoring task, just Mean Squared Error loss was initially used, but later, the combined weighted loss of MSE and Margin Ranking (MR) was implemented. RMSProp[tieleman2012lecture] was used as the optimising algorithm, and scikitlearnscikit-learn’s StandardScalerStandardScaler was used to normalise the scores. The early stoppage technique was used to save computation time if the validation QWK score did not improve in 15 epochs, and validation QWK scores were used to keep the best-performing model.
@@
RMSProp [tieleman2012lecture] was primarily chosen for the AES model for its stable and reliable convergence behaviour. Its gradual adjustments to learning rates make it particularly suited for non-convex optimisation problems, where finding the best solution is challenging. Unlike algorithms like Adam, which can lead to rapid changes in learning rates, RMSProp’s approach avoids potential instability during training, ensuring a smoother learning process.
@@ 5.3 Competing Models
    DeLAES Model: DeLAES, developed by Tashu et al. [tashu2022deep], is a novel approach that employs an n-gram-inspired CNN-Bi-GRU architecture. This model has achieved remarkable success, exhibiting a state-of-the-art mean QWK score of 0.903 across various benchmark datasets. Combining convolutional neural networks (CNNs) and bidirectional gated recurrent units (Bi-GRUs), their innovative strategy to capture local and contextual information in essays for enhanced grading accuracy along the granularities of bi-grams, tri-grams and 4-grams.
@@
    Tran-BERT-MS-ML-R Model: Wang et al. [wang_2022_on] proposed the Tran-BERT-MS-ML-R model, which leverages the power of BERT-based transformer architectures. This model takes a multi-faceted approach by incorporating transfer learning techniques and specialised methodologies to improve the performance of BERT models. By addressing specific challenges associated with essay scoring, Wang et al. achieved notable advancements in scoring accuracy.
@@
    $R^{2}$ BERT Model: $R^{2}$ BERT, introduced by Yang et al. [yang-etal-2020-enhancing], has been the state-of-the-art for transformer-based models for AES for a long time. The name $R^{2}$ BERT stems from its utilisation of both Regression and Ranking techniques for score prediction. This unique combination enhances the model’s ability to capture intricate nuances in essays, improving grading performance. Yang et al. demonstrated the model’s efficacy through comprehensive experimentation and comparisons with existing approaches.
@@
We began by investigating the impact of truncation, a critical consideration for models utilising BERT-based architectures with token limitations. We enhanced the model’s understanding by preserving the direct context of essays through non-truncation. Following this, we further delved into providing a relativistic ordering context via margin ranking, as detailed in Section 4.4. This entailed furnishing the model with a sense of the batch’s essay score ordering to enhance its evaluation capabilities. Further, we examined the augmentation of essays with their respective prompts. This approach improved the model’s grasp of essay quality and its comprehension of nuanced grading patterns within each essay set. We further explored context granularity through the levels of Elementary Discourse Unit (EDU) and Argument Component (AC). Inspired by Wang et al.’s work [wang-etal-2018-toward], we devised models for the upstream tasks of EDU and AC span labelling. These labelled essays formed the input for our transformer-based architecture, where we observed performance improvements.
@@ 6.2 EDU Span Classification
As discussed before, we enhanced the transformer’s grasp of individual text segments and their relationships by leveraging the EDUs. To get EDU labels, we implemented an EDU span classification model, discussed here. Our implementations were based on Wang et al.’s work [wang-etal-2018-toward], and we adapted their architecture to use RoBERTa embeddings instead of Glove and ElMo embeddings.
@@ 6.5 Comparing Models
Tran-BERT-MS-ML-R [wang_2022_on] 0.834 0.716 0.714 0.812 0.813 0.836 0.839 0.766 0.791
$R^{2}$ BERT [yang-etal-2020-enhancing] 0.817 0.719 0.698 0.845 0.841 0.847 0.839 0.744 0.794
@@
  • Graph Convolutional Networks (GCNs): Inspired by the work of Ruiz et al. [ruizdolz_2022_automatic], we may even want to use GCNs to exploit the argument graph structure for further context.
  • Finetuning: Finetuning the transformer-based models is also challenging, as explained in the paper by Wang et al. [wang_2022_on]. We can explore back-translation and relativistic scoring to augment data for further training instances to stabilise fine-tuning as explored by Jong et al. [jong2022improving].
An objective study of these would only positively benefit the literature available in AES. It is important to note that the end-to-end argument graph mining models need to achieve better SOTA results [saha-etal-2022-edu] to avoid injecting noise into the transformer models instead of enriching them.